OpenStudy (anonymous):

One more problem I need help with- if a/4= b/7, then what is the value of (a-b)/b?

OpenStudy (anonymous):

-3/7

Directrix (directrix):

(a-b)/b = (4-7)/7 = -3/7

OpenStudy (anonymous):

Thank you both! c: